OpenStudy (anonymous):

PLEASE HELP. -7m + 20 = -17m - 10

OpenStudy (anonymous):

what i got is m= -3. is that correct?

OpenStudy (rsadhvika):

correct !

OpenStudy (nurali):

-7m + 20 = -17m - 10 -7m+17m=-10-20 10m=-30 m=-30/10 m=-3

OpenStudy (rsadhvika):

you may confirm your work at :- http://www.wolframalpha.com/input/?i=-7m+%2B+20+%3D+-17m+-+10

OpenStudy (anonymous):

thanks!

OpenStudy (nurali):

Anytime.